OpenStudy (anonymous):

Express 108 as a product of primes. Options 108 = 2 x 2 x 3 x 9 108 = 2 x 2 x 3 x 3 x 3 108 = 4 x 27 108 = 2 x 2 x 27

OpenStudy (anonymous):

I think it is the last option

OpenStudy (e.cociuba):

180=2x2x3x3x3

OpenStudy (anonymous):

mmm.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Use prime factorization: http://en.wikipedia.org/wiki/Prime_factor

jimthompson5910 (jim_thompson5910):

it's NOT the last option because 27 is NOT a prime number 27 = 9*3 27 = 3*3*3 which shows you that 27 is NOT prime

OpenStudy (anonymous):

Thank you guys!

OpenStudy (e.cociuba):

yup:)